# thanks: M.C. and D.P. contributed equally to this work

### v1

@@ IV Results
Figure 9: As in Fig. 7, but for primordial regular Ghosh-Culetu-Simpson-Visser black holes, with identical values of the regularizing parameter $\ell/r_{H}$ and identical color coding.
@@
For each of the PRBHs discussed in Sec. II, we now proceed to derive upper limits on $f_{\text{pbh}}$ as a function of the PRBH mass $M_{\text{pbh}}$, for different values of $\ell$, using the methodology presented in Sec. III.3. The results are shown in Figs. 7,  8,  9, and 10 for the Bardeen, Hayward, GCSV, and Hayward-like BHs respectively. For each case, we also plot the constraints on $f_{\text{pbh}}$ for the $\ell=0$ case (blue solid curve in all the Figures), which correspond to the standard Schwarzschild PBH scenario widely studied in the literature. As a sanity check, we have verified that our $\ell=0$ constraints exactly recover those of the seminal Ref. Carr et al. (2010). It is worth noting that, for any given value of $\ell$, the value of $M_{\text{pbh}}$ corresponding to the upper right edge of the $f_{\text{pbh}}$ constraints (i.e. the value of $M_{\text{pbh}}$ for which the limit reads $f_{\text{pbh}}<1$) marks the lower edge of the (modified – either enlarged or contracted) asteroid mass window.
@@
We begin by discussing the Bardeen, Hayward, and GCSV PRBHs, for which we saw earlier that the temperature and photon spectra decrease in intensity with increasing regularizing parameter $\ell$ (see the discussion in Sec. III.2, and Figs. 25). As we could have expected, this behaviour leads to overall looser constraints on $f_{\text{pbh}}$ (for any given $M_{\text{pbh}}$) relative to the standard limits reported for Schwarzschild PBHs in the literature. In the case of near-extremal Hayward PRBHs ($\ell=0.45r_{H}$) this behaviour is somewhat enhanced compared to the near-extremal Bardeen and GCSV PRBHs, with the upper limits on $f_{\text{pbh}}$ approximately three orders of magnitude looser than the corresponding Schwarzschild ones: again, this is somewhat unsurprising when comparing Fig. 3 to Figs. 2 and 4. This could also have been expected from Fig. 1, noting that the temperature of Hayward BHs decreases more rapidly with increasing $\ell/r_{H}$ relative the Bardeen and especially GCSV ones. Although the temperature is not the only factor at play in determining the resulting evaporation spectrum, given that the GBFs also play a key role as per Eq. (29), it is reassuring that the temperature behaviour observed in Fig. 1 is qualitatively reflected in the limits on $f_{\text{pbh}}$ we derive.
@@
For the Hayward-like BH, we observe exactly the opposite trend, with overall tighter constraints on $f_{\text{pbh}}$ relative to the Schwarzschild ones. Again, such a behaviour is in line with expectations given the behaviour of the temperature observed in Fig. 1, and of the spectra as shown in Fig. 5. The relative shift in $f_{\text{pbh}}$ constraints for a given value of $M_{\text{pbh}}$ and $\ell/r_{H}$ is somewhat less dramatic than what we observed for the Bardeen, Hayward, and GCSV BHs, with shifts of less than two orders of magnitude.
@@
As a result of the shifts discussed above, the lower edge of the asteroid mass window where PBHs could make up the entire DM component is modified for all four metrics considered. We recall that in the Schwarzschild case, the lower edge of the window lies at $M_{\text{pbh}}\simeq 10^{14}\,{\text{kg}}$. For the Bardeen, Hayward, and GCSV PRBHs, the looser constraints on $f_{\text{pbh}}$ result in the asteroid mass window further opening up by approximately half a decade in mass or more. The maximum extension of the window is reached for the Hayward PRBH closer to extremality, in which case the lower edge decreases by about an order of magnitude to $M_{\text{pbh}}\simeq 10^{13}\,{\text{kg}}$. The opposite behaviour is of course observed for Hayward-like PRBHs, in which case the asteroid mass window further closes down, although by a lesser extent with respect to the previous cases. For instance, for $\ell/r_{H}=0.75$, the lower edge increases to $M_{\text{pbh}}\simeq 2\times 10^{14}\,{\text{kg}}$. Overall, we therefore observe that considering primordial regular BHs in place of the standard Schwarzschild ones can move the resulting constraints on $f_{\text{pbh}}$ in either direction, further opening up or closing down the asteroid mass window, with the allowed region for the window lower edge spanning over a decade in mass, at least for the PRBHs considered.
@@
Next, the constraints we have determined on $f_{\text{pbh}}$ at a fixed value of $\ell/r_{H}$ implicitly assume that all PRBHs in the Universe carry the same value of “hair” parameter $\ell$. However, particularly given our agnostic stance with regards to the origin of these space-times, in principle the value of $\ell/r_{H}$ can vary from PRBH to PRBH. To make an analogy, let us assume for a moment that Reissner-Nordström BHs are astrophysically relevant. Then, since the electric charge $Q$ is not tied to a universal parameter of the underlying Einstein-Maxwell Lagrangian, there is no reason to expect it to carry the same value across all BHs. In the language of Ref. Vagnozzi et al. (2023), the regularizing parameter for all four RBHs considered is a “specific hair” rather than an “universal hair” (see Ref. Vagnozzi et al. (2023) for various examples of BH solutions carrying universal hair), unless one were able to tie $\ell$ to some fundamental parameter of the underlying theory, which however is not the case in the phenomenological approach we are following. In principle one should therefore account for the (non-monochromatic) $\ell$ distribution for PRBHs across the Universe to determine constraints on $f_{\text{pbh}}$. We see no obvious way of doing this, while noting that such a procedure would most likely result in upper limits on $f_{\text{pbh}}$ lying between the Schwarzschild and extremal cases: this observation suffices for our pilot study, and we defer a more complete investigation to future work.
@@
Of potentially more relevance to the present work are lensing constraints, which constrain the abundance of PBHs (and more generally massive compact halo objects) with masses $M_{\text{pbh}}\gtrsim 10^{12}\,{\text{kg}}$. Indeed, it is lensing constraints which locate the upper edge of the asteroid mass window where PBHs can make up all the DM. Nevertheless, we expect that these constraints should not change when moving from the Schwarzschild PBH framework to the PRBHs considered in this work. Indeed, with all other quantities being fixed (mass of source, relative distances, and so on), lensing constraints only depend on the lens mass $M$, and are unaffected by the metric structure of the lens. Therefore, at fixed mass $M$, we can assume that the lensing limits on Schwarzschild PBHs hold for our PRBHs as well. Note that, as already pointed out in footnote 2, the parameter $M$ appearing in the RBH metrics can be unambiguously identified with the RBH mass, just as with the parameter $M$ in the Schwarzschild metric. We can therefore conclude that for the PRBHs we are considering it is only the lower edge of the asteroid mass window which is altered with respect to the Schwarzschild case, but not the upper edge. In other words, space-times for which the lower edge moves towards lower masses (as in the Bardeen, Hayward, and GCSV PRBH cases) genuinely corresponds to an enlarged asteroid mass window, and conversely for the Hayward-like case where the lower edge moves towards higher masses. Therefore, the window where Bardeen, Hayward, and GCSV PRBHs could account for all the DM is larger compared to that of Schwarzschild PBHs.
Other potentially relevant constraints come from $\mu$-distortions in the Cosmic Microwave Background, and gravitational waves (either a stochastic background due to a population of coalescing PBHs or produced via second-order tensor perturbations generated by the scalar perturbations producing the PBHs, or associated to resolved events). The latter are expected to be relevant in a much higher mass range (again, see Fig. 7 of Ref. Carr and Kuhnel (2022) and Fig. 10 of Ref. Carr et al. (2021d)), whereas the former are somewhat dependent on the PBH formation scenario from high-$\sigma$ tails of density fluctuations, and in particular on the shape of the tail. At any rate, while the focus in the present pilot study has been solely on evaporation constraints from the ERGB, revisiting all these other important sources of constraints (including the ones discussed earlier) is a worthwhile endeavour which we plan to explore in upcoming works.
@@ V Conclusions
Over the past decade, primordial black holes have regained tremendous interest as viable dark matter candidates, with the so-called “asteroid mass window” ($10^{14}\,{\text{kg}}\lesssim M_{\text{pbh}}\lesssim 10^{20}\,{\text{kg}}$) where PBHs could potentially account for the entire DM currently still open. Nearly all works on PBHs assume that these are Schwarzschild or Kerr BHs. However, while phenomenologically perfectly valid, such an assumption may stir some unease on the theoretical side, due to the appearance of singularities in these metrics. In our work, we have conducted a pilot study aimed at addressing a question which naturally merges the DM and singularity problems, arguably two among the most important open problems in theoretical physics: “What if PBHs are non-singular”? Our study of primordial regular BHs (PRBHs) has focused on four so-called tr-symmetric metrics (including the well-known Bardeen and Hayward space-times), whereas our companion paper Calzà et al. (tion) considers non-tr-symmetric metrics, including various metrics inspired from loop quantum gravity.
@@
We show that evaporation constraints on $f_{\text{pbh}}$, the fraction of DM in the form of PRBHs, can be substantially altered in either direction when moving away from the Schwarzschild picture, leading to the asteroid mass window further opening up or closing down depending on the direction of the shifts in $f_{\text{pbh}}$ limits. For three of the PRBHs we considered (the Bardeen, Hayward, and GCSV space-times) the lower edge of the asteroid mass window is shifted by nearly a decade in mass, leading to a larger region of parameter space where PRBHs could account for the entire DM component, which should be the target of the same probes proposed to test the standard window Ray et al. (2021); Ghosh and Mishra (2024); Branco et al. (2023); Bertrand et al. (2023); Tran et al. (2023); Dent et al. (2024); Tamta et al. (2024). The opposite trend takes place with the Hayward-like BH we constructed, and we have argued that part of this different behaviour can be traced back to the evolution (increase or decrease) of the PRBH temperature as the regularizing parameter is increased. On the other hand, the nature of the regular BH core (de Sitter or Minkowski) does not appear to play a significant role in this sense. Overall, we have shown that the phenomenology of primordial regular BHs can be particularly rich, making the associated simultaneous solution to the DM and singularity problems one worthy of further study.
@@
We remark that the present work (alongside our companion paper Calzà et al. (tion)) should be intended as a pilot study, and there are a huge number of interesting follow-up directions. One interesting avenue for further work involves systematically revisiting other sources of constraints which have been studied in the Schwarzschild PBH case, including but not limited to lensing, accretion, and dynamical constraints: while we have argued that these should not alter our considerations on the asteroid mass window, a detailed study which would allow us to extend our constraints over a much larger region of $M_{\text{pbh}}$-$f_{\text{pbh}}$ plane is nevertheless in order. In addition, the metrics we have considered are inherently phenomenological in nature, and it would therefore be worth extending our study to non-singular metrics which enjoy a strong theoretical motivation (our companion paper Calzà et al. (tion) goes partially in this direction), including potentially metrics which are coupled to the cosmological expansion. Last but definitely not least, if PBHs are truly regular, one would hope to ascertain this via signatures complementary to those we have studied (for instance through gravitational wave signatures, VLBI imaging, particle motion, or energy extraction). We plan to address these and other related points in follow-up work.
@@ Acknowledgements.
Appendix A Details on the computation of GBFs
@@
in proximity of its singular points, namely those where $p(x)$ and $q(x)$ diverge. One can notice that Eq. (22) can be rewritten in the form of Eq. (34), with one of its singular point being at $x=0$, i.e. at the event horizon.
@@
$\begin{split}&\tau_{\text{B}}=\frac{1-2\ell^{2}}{\ell^{2}+1}\;\;,\\ &\tau_{\text{H}}=(1-3\ell^{2})\;\;,\\ &\tau_{\text{GCSV}}=1-\ell\;\;,\\ &\tau_{\text{H-l}}=\frac{(\ell-1)^{2}}{1-\ell(\ell-1)}\;\;.\end{split}$

### v2

@@ IV Results
Figure 9: As in Fig. 7, but for primordial regular Culetu-Ghosh-Simpson-Visser black holes, with identical values of the regularizing parameter $\ell/r_{H}$ and identical color coding.
@@
For each of the PRBHs discussed in Sec. II, we now proceed to derive upper limits on $f_{\text{pbh}}$ as a function of the PRBH mass $M_{\text{pbh}}$, for different values of $\ell$, using the methodology presented in Sec. III.3. The results are shown in Figs. 7,  8,  9, and 10 for the Bardeen, Hayward, CGSV, and Hayward-like PRBHs respectively. For each case, we also plot the constraints on $f_{\text{pbh}}$ for the $\ell=0$ case (blue solid curve in all the Figures), which correspond to the standard Schwarzschild PBH scenario widely studied in the literature. As a sanity check, we have verified that our $\ell=0$ constraints exactly recover those of the seminal Ref. Carr et al. (2010). It is worth noting that, for any given value of $\ell$, the value of $M_{\text{pbh}}$ corresponding to the upper right edge of the $f_{\text{pbh}}$ constraints (i.e. the value of $M_{\text{pbh}}$ for which the limit reads $f_{\text{pbh}}<1$) marks the lower edge of the (modified – either enlarged or contracted) asteroid mass window.
@@
We begin by discussing the Bardeen, Hayward, and CGSV PRBHs, for which we saw earlier that the temperature and photon spectra decrease in intensity with increasing regularizing parameter $\ell$ (see the discussion in Sec. III.2, and Figs. 25). As we could have expected, this behaviour leads to overall looser constraints on $f_{\text{pbh}}$ (for any given $M_{\text{pbh}}$) relative to the standard limits reported for Schwarzschild PBHs in the literature. In the case of near-extremal Hayward PRBHs ($\ell=0.45r_{H}$) this behaviour is somewhat enhanced compared to the near-extremal Bardeen and CGSV PRBHs, with the upper limits on $f_{\text{pbh}}$ approximately three orders of magnitude looser than the corresponding Schwarzschild ones: again, this is somewhat unsurprising when comparing Fig. 3 to Figs. 2 and 4. This could also have been expected from Fig. 1, noting that the temperature of Hayward BHs decreases more rapidly with increasing $\ell/r_{H}$ relative to the Bardeen and especially CGSV ones. Although the temperature is not the only factor at play in determining the resulting evaporation spectra, given that the GBFs also play a key role as per Eq. (29), it is reassuring to see that the temperature behaviour observed in Fig. 1 is qualitatively reflected in the limits on $f_{\text{pbh}}$ we derive.
@@
For the Hayward-like BH, we observe exactly the opposite trend, with overall tighter constraints on $f_{\text{pbh}}$ relative to the Schwarzschild ones. Again, such a behaviour is in line with expectations given the behaviour of the temperature observed in Fig. 1, and of the spectra as shown in Fig. 5. The relative shift in $f_{\text{pbh}}$ constraints for a given value of $M_{\text{pbh}}$ and $\ell/r_{H}$ is somewhat less dramatic than what we observed for the Bardeen, Hayward, and CGSV BHs, with shifts of less than two orders of magnitude.
@@
As a result of the shifts discussed above, the lower edge of the asteroid mass window where PBHs could make up the entire DM component is modified for all four metrics considered. We recall that in the Schwarzschild case, the lower edge of the window lies at $M_{\text{pbh}}\simeq 10^{14}\,{\text{kg}}$. For the Bardeen, Hayward, and CGSV PRBHs, the looser constraints on $f_{\text{pbh}}$ result in the asteroid mass window further opening up by up to half a decade in mass or more. The maximum extension of the window is reached for the Hayward PRBH closer to extremality, in which case the lower edge decreases by about an order of magnitude to $M_{\text{pbh}}\simeq 10^{13}\,{\text{kg}}$. The opposite behaviour is of course observed for Hayward-like PRBHs, in which case the asteroid mass window further closes down, although by a lesser extent with respect to the previous cases. For instance, for $\ell/r_{H}=0.75$, the lower edge increases to $M_{\text{pbh}}\simeq 2\times 10^{14}\,{\text{kg}}$. Overall, we therefore observe that considering PRBHs in place of the standard Schwarzschild ones can move the resulting constraints on $f_{\text{pbh}}$ in either direction, further opening up or closing down the asteroid mass window, with the allowed region for the window lower edge spanning over a decade in mass, at least for the PRBHs considered.
@@
Next, the constraints we have determined on $f_{\text{pbh}}$ at a fixed value of $\ell/r_{H}$ implicitly assume that all PRBHs in the Universe carry the same value of “hair” parameter $\ell$. However, particularly given our agnostic stance with regards to the origin of these space-times, in principle the value of $\ell/r_{H}$ can vary from PRBH to PRBH. To make an analogy, let us assume for a moment that Reissner-Nordström BHs are astrophysically relevant. Then, since the electric charge $Q$ is not tied to a universal parameter of the underlying Einstein-Maxwell Lagrangian, there is no reason to expect it to carry the same value across all BHs. In the language of Ref. Vagnozzi et al. (2023), the regularizing parameter for all four PRBHs considered is a “specific hair” rather than an “universal hair” (see Ref. Vagnozzi et al. (2023) for various examples of BH solutions carrying universal hair), unless one were able to tie $\ell$ to some fundamental parameter of the underlying theory, which however is not the case in the phenomenological approach we are following. In principle one should therefore account for the (non-monochromatic) $\ell$ distribution for PRBHs across the Universe to determine constraints on $f_{\text{pbh}}$. We see no obvious way of doing this, while noting that such a procedure would most likely result in upper limits on $f_{\text{pbh}}$ lying between the Schwarzschild and extremal cases: this observation suffices for our pilot study, and we defer a more complete investigation to future work.
@@
Of potentially more relevance to the present work are lensing constraints, which constrain the abundance of PBHs (and more generally massive compact halo objects) with masses $M_{\text{pbh}}\gtrsim 10^{12}\,{\text{kg}}$. Indeed, it is lensing constraints which locate the upper edge of the asteroid mass window where PBHs can make up all the DM. Nevertheless, we expect that these constraints should not change when moving from the Schwarzschild PBH framework to the PRBHs considered in this work. Indeed, with all other quantities being fixed (mass of source, relative distances, and so on), lensing constraints only depend on the lens mass $M$, and are unaffected by the metric structure of the lens. Therefore, at fixed mass $M$, we can assume that the lensing limits on Schwarzschild PBHs hold for our PRBHs as well. Note that, as already pointed out in footnote 2, the parameter $M$ appearing in the RBH metrics can be unambiguously identified with the RBH mass, just as with the parameter $M$ in the Schwarzschild metric. We can therefore conclude that for the PRBHs we are considering it should only be the lower edge of the asteroid mass window which is altered with respect to the Schwarzschild case, but not the upper edge. In other words, space-times for which the lower edge moves towards lower masses (as in the Bardeen, Hayward, and CGSV PRBH cases) genuinely correspond to an enlarged asteroid mass window, and conversely for the Hayward-like case where the lower edge moves towards higher masses. Therefore, the window where Bardeen, Hayward, and CGSV PRBHs could account for all the DM is much larger compared to that of Schwarzschild PBHs.
Other potentially relevant constraints come from $\mu$-distortions in the Cosmic Microwave Background, and gravitational waves (either a stochastic background due to a population of coalescing PBHs or produced via second-order tensor perturbations generated by the scalar perturbations producing the PBHs, or associated to resolved events). The latter are expected to be relevant in a much higher mass range (again, see Fig. 7 of Ref. Carr and Kuhnel (2022) and Fig. 10 of Ref. Carr et al. (2021d)), whereas the former are somewhat dependent on the PBH formation scenario from high-$\sigma$ tails of density fluctuations, and in particular on the shape of the tail. At any rate, while the focus in the present pilot study has been solely on evaporation constraints from the ERGB, revisiting all these other important sources of constraints (including the ones we discussed earlier) is a worthwhile endeavour which we plan to explore in upcoming works.
@@ V Conclusions
Over the past decade, primordial black holes have regained tremendous interest as viable dark matter candidates, with the so-called “asteroid mass window” ($10^{14}\,{\text{kg}}\lesssim M_{\text{pbh}}\lesssim 10^{20}\,{\text{kg}}$) where PBHs could potentially account for the entire DM currently still open. Nearly all works on PBHs assume that these are Schwarzschild or Kerr BHs. However, while phenomenologically perfectly valid, such an assumption may stir some unease on the theoretical side, due to the appearance of singularities in these metrics. In our work, we have conducted a pilot study aimed at addressing a question which naturally merges the DM and singularity problems, arguably two among the most important open problems in theoretical physics: “What if PBHs are non-singular”? Our study of primordial regular BHs (PRBHs) has focused on four so-called tr-symmetric metrics (including the well-known Bardeen and Hayward space-times), whereas our companion paper Calzà et al. (2024a) considers non-tr-symmetric metrics, including various metrics inspired from loop quantum gravity.
@@
We show that evaporation constraints on $f_{\text{pbh}}$, the fraction of DM in the form of PRBHs, can be substantially altered in either direction when moving away from the Schwarzschild picture, leading to the asteroid mass window further opening up or closing down depending on the direction of the shifts in $f_{\text{pbh}}$ limits. For three of the PRBHs we considered (the Bardeen, Hayward, and CGSV ones) the lower edge of the asteroid mass window is shifted by nearly a decade in mass, leading to a larger region of parameter space where PRBHs could account for the entire DM component, which should be the target of the same probes proposed to test the standard window Ray et al. (2021); Ghosh and Mishra (2024); Branco et al. (2023); Bertrand et al. (2023); Tran et al. (2023); Dent et al. (2024); Tamta et al. (2024). The opposite trend takes place with the Hayward-like BH we constructed, and we have argued that part of this different behaviour can be traced back to the evolution (increase or decrease) of the PRBH temperature as the regularizing parameter is increased. On the other hand, the nature of the regular BH core (de Sitter or Minkowski) does not appear to play a significant role in this sense. Overall, we have shown that the phenomenology of primordial regular BHs can be particularly rich, making the associated simultaneous solution to the DM and singularity problems one worthy of further study.
@@
We remark that the present work (alongside our companion paper Calzà et al. (2024a)) should be intended as a pilot study, and there are a huge number of interesting follow-up directions. One interesting avenue for further work involves systematically revisiting other sources of constraints which have been studied in the Schwarzschild PBH case, including but not limited to lensing, accretion, and dynamical constraints: while we have argued that these should not alter our considerations on the asteroid mass window, a detailed study which would allow us to extend our constraints over a much larger region of $M_{\text{pbh}}$-$f_{\text{pbh}}$ plane is nevertheless in order. In addition, the metrics we have considered are inherently phenomenological in nature, and it would therefore be worth extending our study to non-singular metrics which enjoy a strong theoretical motivation (our companion paper Calzà et al. (2024a) goes partially in this direction), including potentially metrics which are coupled to the cosmological expansion Farrah et al. (2023); Cadoni et al. (2023, 2024); Faraoni and Rinaldi (2024); Calzà et al. (2024b). Last but definitely not least, if PBHs are truly regular, one would hope to ascertain this via signatures complementary to those we have studied (for instance through gravitational wave signatures, VLBI imaging, particle motion, or energy extraction). We plan to address these and other related points in follow-up work.
@@ Acknowledgements.
Appendix A Details on the computation of greybody factors
@@
in proximity of its singular points, namely those where $p(x)$ and/or $q(x)$ diverge. One can notice that Eq. (22) can be rewritten in the form of Eq. (34), with one of its singular point being at $x=0$, i.e. at the event horizon.
@@
$\begin{split}&\tau_{\text{B}}=\frac{1-2\ell^{2}}{\ell^{2}+1}\;\;,\\ &\tau_{\text{H}}=(1-3\ell^{2})\;\;,\\ &\tau_{\text{CGSV}}=1-\ell\;\;,\\ &\tau_{\text{H-l}}=\frac{(\ell-1)^{2}}{1-\ell(\ell-1)}\;\;.\end{split}$